OpenStudy (anonymous):

Use the intermediate value theorem to show that there is a root (solution) of the given equation in the interval (1,2) x^3-2x=3

OpenStudy (anonymous):

put \(f(x)=x^3-2x\) and compute \(f(1)\) and \(f(2)\) one is since \(f(1)<3\) and \(f(2)>3\) there must be a number \(c\) in the interval (1,2) where \(f(c)=3\)

OpenStudy (anonymous):

k i would get f(1)=3 and f(2)=0

OpenStudy (anonymous):

Well 1<2 a<b f(x)=x^3-2x-3=0 You have to plug 1 and 2 into the equation (1)^3-2(1)-3 1-2-3 1-5 -4 (2)^3-2(2)-3 8-4-3 8-7 1 Because f(1)<0 and f(2)>0 and f is continuous, therefore there exist a zero in the interval 1 and 2.

OpenStudy (anonymous):

minus 3 from both sides of the equation

OpenStudy (anonymous):

k didn't c that